shyane:

Sandy is working with a carpenter to frame a house. They are using 8-foot-long boards, but each board must be cut to be 94.6 inches long. How much is cut off each board? (1 foot = 12 inches) meh need sum help

dude:

Okay so you have to convert 8 feet to inches They gave you a ratio (1 ft = 12 in) Use proportions to find the measure in inches first \(\large\frac{8ft}{}\cdot\frac{12in}{1ft}\)

dude:

From then subtract the amount used per board (94.6)